OpenStudy (anonymous):

just looking to get my answer checked! 6e^x/sinx for the derivative i have: -6 g(x)=sinx-cosx/sin^2(x)

OpenStudy (turingtest):

how did the 6 get over there? and what is g(x) and where did e^x go?

OpenStudy (anonymous):

i thought they would cancel but, i just looked at my workings and that doesn't make sense because it's sinx(6e^x)-(6e^x)(cosx) g(x) is stated above in my question and i pulled out the 6 because it's a constant

OpenStudy (turingtest):

so g(x)=6e^x/sinx ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and i need to find the derivative.

OpenStudy (turingtest):

well you have the right numerator so far sinx(6e^x)-(6e^x)(cosx) and the denominator?

OpenStudy (anonymous):

sin2x

OpenStudy (anonymous):

thats sin squared.

OpenStudy (anonymous):

so i can factor out 6e^x on the top.

OpenStudy (turingtest):

exactly \[g'(x)=\frac{6e^x\sin x-6e^x\cos x}{\sin^2x}\]

OpenStudy (anonymous):

thank youuuu!

OpenStudy (turingtest):

welcome!